Ayie is 8 yrs. old than Icas. The sum of their age is 32. How old is each now?
Found 2 solutions by rapaljer, meay7c:
Answer by rapaljer(4671)

Let x=icas age
x+8= ayie age

x+x+8=32
2x+8=32
2x=24
x=12 icas
x+8=20 ayie

The sum of their ages is 32 so it checks!!

For additional help on WORD PROBLEMS, please see my own website for a LOT of FREE resources. Click on my tutorname "rapaljer" anywhere in algebra.com. Once you find my Homepage, look for "Basic, Intermediate, and College Algebra: One Step at a Time". There you will find my own user-friendly explanation for just about all the topics in algebra. Choose "Basic Algebra", then look in "Chapter 1" for "Sections 1.09 and 1.10".

Many of my exercises in this section are solved in the "MATH IN LIVING COLOR" sections that go with this. There you will find an explanation that will probably be a lot easier to understand than your own expensive textbooks!! My materials are all free.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus


Answer by meay7c(10)

Let x = Icas age
We know that Ayie's age will be 8 more than Icas' age. (8 + x)
The sum of both ages is 32.
Let's set up an equation:
(8 + x) + x = 32
= 8 + (x + x) = 32
= 8 + 2x =32
= 2x = 24
x = 12
So Icas is 12 and
Ayie is 20 (8 + 12 = 20)
Check your solutions by plugging x into the first equation:
8 + 12 + 12 = 32 Yes!!!